OpenStudy (anonymous):

The sum of the interior angles of a regular polygon is one thousand eight hundred degrees. What is the measure of each exterior angle of the polygon? One hundred fifty degrees One hundred eight degrees Sixty degrees Thirty degrees

OpenStudy (cwrw238):

hint sum of exterior angles = 360 degrees

OpenStudy (anonymous):

I know that already but I'm still confused :(

OpenStudy (mathstudent55):

In a regular polygon, all interior angles are congruent. Also, the sum of the measures of the interior angles of a polygon of n sides is 180(n - 2) 

OpenStudy (mathstudent55):

In the figure above, x represents the measure of one interior angle. y is the measure of one exterior angle.

OpenStudy (anonymous):

I learned all this information from my lesson but I just don't know the answer :(

OpenStudy (mathstudent55):

Look at the formula of the sum of the measures of the interior angles. S = 180(n - 2) You know S, the sum. If you plug it in, you can find the number of the sides, n.

OpenStudy (cwrw238):

if exterior angle = x n = no.of sides nx = 360 180(n - 2) = 1080

OpenStudy (mathstudent55):

180(n - 2) = 1800 Solve for n to find the number of sides in the polygon.

OpenStudy (anonymous):

how do I solve for n?

OpenStudy (mathstudent55):

180(n - 2) = 1800 First, divide both sides by 180.

OpenStudy (anonymous):

ok I did that now what?

OpenStudy (mathstudent55):